# Functional Homotopy: Smoothing Discrete Optimization via Continuous Parameters for LLM Jailbreak Attacks

### v1

@@ 1 Introduction
Optimization techniques for generating malicious inputs have been extensively applied in adversarial learning, particularly within image models. The most prevalent methods include gradient-based approaches such as the Fast Gradient Sign Method (FGSM) [9] and Projected Gradient Descent (PGD) [15]. These techniques have demonstrated that many deep learning models exhibit vulnerability to small $\ell_{p}$ perturbations to the input. The optimization problem for generating malicious inputs can be expressed as:
@@
For language models, researchers have also utilized optimization techniques to generate inputs that provoke extreme undesired behaviors. Approaches analogous to those employed in adversarial learning have been adopted for this purpose. For example, Greedy Coordinate Gradient [25] (GCG) employs gradient-based methods to identify tokens that induce jailbreak behaviors. Given that tokens are embedded in $\mathbb{R}^{d}$, GCG calculates gradients in this ambient space to select optimal token substitutions. This methodology has also been adopted by other studies for related prompt synthesis challenges [11, 14].
@@
In this paper, we introduce a novel optimization method for addressing the problem formulated in Equation 1, specifically when the input variable $x$ resides in a discrete space. Direct optimization of this problem within $\mathcal{X}\subset\mathbb{R}^{n}$ using token gradient methods is insufficient, as gradients provide only local information, which often fails to account for the substantial distances between tokens in the ambient space. However, although combinatorial optimization problems are generally classified as $\mathsf{NP}$-hard [21], the problem in Equation 1 exhibits a unique characteristic: the function $f_{p}$ is parameterized by $p$, which lies in a continuous domain. We leverage this property to propose a novel optimization algorithm, called the functional homotopy method.
The homotopy method [7] involves gradually transforming a challenging optimization problem into a sequence of easier problems, utilizing the solution from the previous problem to warm start the optimization process of the next problem. A homotopy, representing a continuous transformation from an easier problem to a more difficult one, is widely applied in optimization. For instance, the well-known interior point method for constrained optimization by constructing a series of soft-to-hard constraints [2]. Various approaches exist for constructing a homotopy, such as employing parameterized penalty terms, as demonstrated in the interior point method, or incorporating Gaussian random noise [18].
In our functional homotopy (FH) method, we go beyond the conventional interpretation of $f_{p}$ in Equation 1 as a static objective function, which was the perspective taken in previous work [25, 13, 11, 1]. Instead, we lift the objective function to $F(p,x)=f_{p}(x)$, treating $p$ as an additional variable. Equation 1 thus becomes:
@@
In the FH method for Equation 2, we first optimize over the continuous parameter $p$. Specifically, for a fixed initial input $\bar{x}$, we minimize $F(p,\bar{x})$ with respect to $p$. We employ gradient descent to update $p$ until a desired value of $F(p^{\prime},\bar{x})$ is achieved. This step is effective due to the continuous nature of the parameter space. As the parameter $p$ is iteratively updated in this process, we can retain all intermediate states of the parameter, denoted as $p_{0}=p,p_{1},\ldots,p_{t}=p^{\prime}$.
@@
  • We present a quantitative analysis of the effectiveness of token gradients on the underlying optimization problem (see Section 4) and characterize its potential efficacy, which depends on the accuracy of the linear approximation of the objective function. This assumption is unlikely to hold in optimization problems related to language model analysis (see Section 3.2).
@@
  • Our application of this algorithm to jailbreak attack generation shows that our method surpasses existing optimization techniques, achieving a 20% to 30% improvement in success rate when circumventing established safe open-source models (see Section 5).
@@ 2 Related Work
Research has demonstrated that neural networks in image models are particularly susceptible to adversarial attacks generated through optimization techniques [19, 3]. In response, researchers have developed adversarially robust models using a min-max saddle-point formulation [15]. Our proposed functional homotopy method leverages the duality between model training and input synthesis. Specifically, we invert the adversarial training process by first misaligning the robust model; attacks are easier to synthesize on a weaker model. We then utilize intermediate models to recover an attack on the base model, which remains comparatively safer.
@@
In recent years, there has been a significant increase in interest regarding jailbreak attacks on LLMs. Various methodologies have been explored, including manual red teaming efforts [8, 20, 22, 1], leveraging other LLMs to compromise target models [17, 4], and automating jailbreak generation through optimization techniques [25, 13, 11]. Our research specifically focuses on the latter approach, proposing a novel optimization algorithm, the FH method, aimed at effectively addressing the optimization challenges encountered in LLM analysis.
@@ 3 Method
In this section, we reevaluate the token gradient method, demonstrating its limitations in effectively addressing the underlying optimization problem. Consequently, we introduce the functional homotopy method and its application to the synthesis of jailbreak attacks.
@@ 3.1 Notations and definitions
    Given a prompt $p$, the output of $M$, denoted by $M(p)\in\Delta(V^{*})$, is a probability distribution over token sequences. $\Delta(V^{*})$ denotes the probability simplex on $V^{*}$.
@@
    Let $T(M(p))\in V^{*}$ be the realized output answer of $M$ to the prompt $p$, where the tokens of $T(M(p))$ are drawn from the distribution $M(p)$.
@@
Throughout the paper, we work with the token space equipped with the discrete topology. We often refer to $\mathcal{X}$ as a topological space when the context is unambiguous.
@@
3.2 Token Gradient Methods
In this section we revisits existing gradient methodologies applied to the token space $\mathcal{X}$, highlighting that their effectiveness hinges on the accuracy of the linear approximation of the objective in Equation 1. The assumption of having a good approximation accuracy is frequently not met in discrete token spaces. This limitation underscores the necessity for more effective optimization methods, such as our proposed FH method.
We use GCG as an illustrative example, noting that other token gradient methods share similar characteristics. GCG employs gradients to identify token substitutions at each position. For an input $x_{0}$, we compute the gradient of $f$ at $x_{0}$, denoted as $Df(x_{0})$. The gradient $Df(x_{0})$ has the same dimensionality as $x_{0}$. At position $j$, let $h=Df(x_{0})_{j}\in\mathbb{R}^{n}$ be the $j$-th component of $Df(x_{0})$. We can compute $k=\operatorname*{arg\,max}(h)$, which corresponds to the $k$-th token in the vocabulary $V$. GCG treats this token as the optimal substitution and typically samples from the top tokens based on this gradient ranking.
@@ 3.2 Token Gradient Methods
The proof is presented in Section A.1.1. Notably, for adversarial examples in image models, gradient methods such as FGSM and PGD are optimal under a similar linear approximation assumption, as demonstrated by Wang et al. [23]. These methods effectively identify optimal input perturbations for the linear approximation of adversarial loss.
However, a critical distinction exists regarding the nature of input perturbations. In adversarial examples, perturbations are confined to small continuous $\ell_{p}$-balls, facilitating precise linear approximations. Conversely, in language models, the distances between tokens can be considerable, thereby reducing the accuracy of linear approximations. Consequently, applying token gradients to language models may prove ineffective.
@@ 3.3 Functional Homotopy method
The optimization problem $\min_{x\in\mathcal{X}}f(x)$ is a nonconvex and hard problem, whereas $\min_{x\in\mathcal{X}}g(x)$ is an easy optimization problem. As a result, $H(x,t)$ induces a series of easy-to-hard optimization problems as $t$ goes from 00 to $1$.
One can then gradually solve this series of problems, by warm starting the optimization algorithm using the solution from the previous similar problem and eventually solve $\min_{x\in\mathcal{X}}f(x)$. Figure 2 illustrates an example of homotopy from $g(x)$ to $f(x)$. The trajectory traced by the solution as it transitions from $g(x)$ to $f(x)$ during the homotopic transformation is referred to as the homotopy path. Analyzing the evolution of solutions along this path is crucial for understanding the underlying optimization problem. For instance, in the interior point method, the homotopy path evolution provides the convergence analysis of the algorithm [2].
@@
Constructing a homotopy offers various approaches, such as utilizing parameterized penalty terms (as in the interior point method) or incorporating Gaussian random noise [18]. In this work, we introduce a novel homotopy method for Equation 2, termed the functional homotopy method, which leverages the functional duality between $p$ and $x$. Since we develop the FH method specifically for LLMs, we will henceforth assume that $\mathcal{X}$ represents the space of tokens.
@@
By allowing infinitesimal updates (learning rates), the gradient descent over the parameter space creates a homotopy between $F(p,x)$ and $F(p^{\prime},x)$, with $H(x,t=0)=F(p^{\prime},x)$ and $H(x,t=1)=F(p,x)$ for the homotopy method. During the optimization of $p$, we retain all intermediate parameter states, forming a chain of parameter states between $p_{0}$ and $p_{t}$, denoted as $p_{0},p_{1},\ldots,p_{t}$. Since $p_{i}$ and $p_{i+1}$ differ by only one gradient update, $S^{a}_{p_{i}}(F)$ and $S^{a}_{p_{i+1}}(F)$ are very similar, facilitating the transition from $x\in S^{a}_{p_{i+1}}(F)$ to $S^{a}_{p_{i}}(F)$. A formal description of the functional homotopy algorithm is provided in Algorithm 1.
@@
Machine Learning Interpretation
This algorithm can be interpreted through a machine learning lens. In the first step, the algorithm updates $p$ with respect to the objective function $F(p,x_{t})$, which can be viewed as training or fine-tuning the model. When $F$ quantifies the safety of a model, this update can be considered as misaligning the model. By preserving the intermediate states of $p$, we obtain a sequence of models ranging from weak to strong safety. Subsequently, we attack this model sequence, initializing each attack from the preceding successful attack on a weaker model. Given that $p_{i}$ and $p_{i+1}$ are closely related, the attacks on these models are also likely to be similar. Thus, we effectively decompose a complex optimization problem into a series of progressively challenging subproblems.
3.4 Application
@@
For jailbreak attack generation, the objective is to persuade $M$ to provide an unaligned and potentially harmful response to a malicious query $x$ (e.g., “how to make a bomb?”), rather than refusing to answer. If $M$ is well-aligned, $T(M(p))$ should result in a refusal. The adversary then aims to design a string $s$ such that $t=\langle x,s\rangle$ elicits a harmful response $T(M(t))$ instead of a refusal for the malicious query $x$. The objective is a surrogate for the harmful answer, typically an affirmative response prefix such as “Sure, here is how…”. Zou et al. [25], Liu et al. [13], Hu et al. [11] have adopted similar formalizations for jailbreak generation.
@@ 4.1 Experimental Design
To quantify the similarity between these rankings, we employ the Rank Biased Overlap (RBO) metric [24]. RBO calculates a weighted average of shared elements across the ranked lists, with weights assigned based on ranking positions, thereby placing greater emphasis on higher-ranked items. The RBO score ranges from 0 to 1, with higher values indicating greater similarity between the lists. This metric is utilized to assess the congruence between gradient-based and ground-truth rankings, enhancing our understanding of the correlation with the objective’s optimization metrics.
@@
We conduct a similar experiment to RQ2, but we record the number of search iterations used by each tool.
@@ 4.2 Experimental Specifications
GR operates as a token-level search algorithm similar to GCG; however, it uses random selection for token substitutions rather than utilizing gradient information. This algorithm serves as an end-to-end implementation of 3 within Algorithm 1. Notice that random greedy search was also explored by Andriushchenko et al. [1], as part of a bag of tricks applied in the work. Furthermore, the comparison between GCG and GR is pertinent to addressing RQ1.
@@
We use recent open source state-of-the-art models, in terms of performance and robustness. These include: Llama-3 8B Instruct [6], Llama-2 7B [20], Mistral-v0.3 7B Instruct [12] and Vicuna-v1.5 7B [5].
@@
For RQ1, we select 20 samples from the AdvBench dataset [25] and randomly choose four positions in the suffix for token substitution for each sample. For each query and position, we substitute all possible tokens (32 0003200032\,00032 000 for Llama-2, Mistral, and Vicuba; 128 256128256128\,256128 256 for Llama-3) and evaluate the jailbreak loss values using these inputs as ground truth, thereby establishing a ground truth ranking. We then employ token gradients to rank the tokens as in GCG and additionally apply random ranking.
For RQ2 and RQ3, we utilize 100 random samples from both the AdvBench and HarmBench datasets [16], resulting in a total of 200 samples. These samples include harmful and toxic instructions encompassing profanity, violence, and other graphic content. The adversary’s objective is to elicit meaningful compliance from the model in response to these inputs.
@@
We utilize the Llama-2 13B model, as provided by [16], to evaluate the responses generated through adversarial attacks, specifically measuring the success rate of these attacks. In the context of jailbreak attack synthesis, the primary objective is to pass the evaluation by the judge, which effectively corresponds to the set $S_{p}^{a}(F)$ in Algorithm 1.
@@
The initial step of our FH method involves updating $p$, which effectively corresponds to model fine-tuning. To optimize memory and disk efficiency while preserving all intermediate parameter states, we employ Low-Rank Adaptation (LoRA) [10] for updating $p$. Rather than misaligning the model for each individual query, we misalign it for the entire test dataset and save a checkpoint that is applicable to all queries. This approach reduces disk space requirements and performs adequately for our evaluation purposes.
@@
We include other experimental specifications in the appendix.

### v2

@@ 1 Introduction
Optimization techniques for generating malicious inputs have been extensively applied in adversarial learning, particularly to image models. The most prevalent methods include gradient-based approaches such as the Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015) and Projected Gradient Descent (PGD) (Madry et al., 2018). These techniques have demonstrated that many deep learning models exhibit vulnerability to small $\ell_{p}$ perturbations to the input. The optimization problem for generating malicious inputs can be expressed as:
@@
For language models, researchers have also utilized optimization techniques to generate inputs that provoke extreme undesired behaviors. Approaches analogous to those employed in adversarial learning have been adopted for this purpose. For example, Greedy Coordinate Gradient (Zou et al., 2023) (GCG) employs gradient-based methods to identify tokens that induce jailbreak behaviors. Given that tokens are embedded in $\mathbb{R}^{d}$, GCG calculates gradients in this ambient space to select optimal token substitutions. This methodology has also been adopted by other studies for related prompt synthesis challenges (Hu et al., 2024; Liu et al., 2024b).
@@
In this paper, we establish that the model-agnostic optimization problem for LLM input generation is $\mathsf{NP}$-hard, implying that efficient optimization algorithms likely require exploitation of problem-specific structures. To address this challenge, we propose a novel optimization method tailored to the problem defined in Equation 1. This approach can be interpreted as utilizing the model-alignment property by intentionally modifying the model to reduce its alignment. Despite the discrete input space, the problem in Equation 1 exhibits a unique characteristic: the function $f_{p}$ is parameterized by $p$, which lies in a continuous domain. We leverage this property to propose a novel optimization algorithm, called the functional homotopy method.
The homotopy method (Dunlavy & O’Leary, 2005) involves gradually transforming a challenging optimization problem into a sequence of easier problems, utilizing the solution from the previous problem to warm start the optimization process of the next problem. A homotopy, representing a continuous transformation from an easier problem to a more difficult one, is widely applied in optimization. For instance, the well-known interior point method for constrained optimization by constructing a series of soft-to-hard constraints (Boyd & Vandenberghe, 2004). Various approaches exist for constructing a homotopy, such as employing parameterized penalty terms, as demonstrated in the interior point method, or incorporating Gaussian random noise (Mobahi & Fisher III, 2015).
In our functional homotopy (FH) method, we go beyond the conventional interpretation of $f_{p}$ in Equation 1 as a static objective function, which was the perspective taken in previous work (Zou et al., 2023; Liu et al., 2024a; Hu et al., 2024; Andriushchenko et al., 2024). Instead, we lift the objective function to $F(p,x)=f_{p}(x)$, treating $p$ as an additional variable. Equation 1 thus becomes:
@@
In the FH method for Equation 2, we first optimize over the continuous parameter $p$. Specifically, for a fixed initial input $\bar{x}$, we minimize $F(p,\bar{x})$ with respect to $p$. We employ gradient descent to update $p$ until a desired value of $F(p^{\prime},\bar{x})$ is achieved. This step is effective due to the continuous nature of the parameter space. As the parameter $p$ is iteratively updated in this process, we retain all intermediate states of the parameter, denoted as $p_{0}=p,p_{1},\ldots,p_{t}=p^{\prime}$.
@@
  • We establish that the model-agnostic optimization problem for LLM input generation is $\mathsf{NP}$-hard (see Section 3.2).
@@
  • Our application of this algorithm to jailbreak attack generation shows that our method surpasses existing optimization techniques, achieving a 20% to 30% improvement in success rate when circumventing established safe open-source models (see Section 5). Our findings also reveal that intermediate model checkpoints can facilitate attacks on base models, offering significant insights for the deep learning community. The evaluation repository is available at: https://github.com/danshumaan/functional_homotopy.
@@ 2 Related Work
Research has demonstrated that neural networks in image models are particularly susceptible to adversarial attacks generated through optimization techniques (Szegedy et al., 2014; Carlini & Wagner, 2017). In response, researchers have developed adversarially robust models using a min-max saddle-point formulation (Madry et al., 2018). Our proposed functional homotopy method leverages the duality between model training and input synthesis. Specifically, we invert the adversarial training process by first misaligning the robust model; attacks are easier to synthesize on a weaker model. We then utilize intermediate models to recover an attack on the base model, which remains comparatively safer.
@@
In recent years, there has been a significant increase in interest regarding jailbreak attacks on LLMs. Various methodologies have been explored, including manual red teaming efforts (Ganguli et al., 2022; Touvron et al., 2023; walkerspider, 2022; Andriushchenko et al., 2024), leveraging other LLMs to compromise target models (Mehrotra et al., 2023; Chao et al., 2024), and automating jailbreak generation through optimization techniques (Schwinn et al., 2024; Zou et al., 2023; Liu et al., 2024a; Hu et al., 2024; Liao & Sun, 2024). Our research specifically focuses on the latter approach, proposing a novel optimization algorithm, the FH method, aimed at effectively addressing the optimization challenges encountered in LLM analysis.
@@ 3 Method
This section presents the functional homotopy method and its application to jailbreak attack synthesis. Additionally, we provide a proof demonstrating that the model-agnostic optimization problem for jailbreak synthesis is $\mathsf{NP}$-hard.
@@ 3.1 Notations and definitions
    Given a prompt $x$, the output of $M$, denoted by $M(x)\in\Delta(V^{*})$, is a probability distribution over token sequences. $\Delta(V^{*})$ denotes the probability simplex on $V^{*}$.
@@
    Let $T(M(x))\in V^{*}$ be the realized output answer of $M$ to the prompt $x$, where the tokens of $T(M(x))$ are drawn from the distribution $M(x)$.
@@
Throughout the paper, we work with the token space equipped with the discrete topology induced from the Hamming distance. We often refer to $\mathcal{X}$ as a topological space when the context is clear.
@@
3.2 Hardness of Jailbreak Attack synthesis
In this section, we establish that the model-agnostic LLM input generation optimization problem is $\mathsf{NP}$-hard. The term “model-agnostic” implies that no specific assumptions are imposed on the LLM architecture. Additionally, we analyze existing gradient-based methods applied to the token space $\mathcal{X}$, emphasizing their reliance on the accuracy of the linear approximation of the objective function in Equation 1. However, this assumption often fails in discrete token spaces, underscoring the necessity for more robust optimization techniques.
Theorem 3.1.
The model-agnostic LLM input generation optimization problem is $\mathsf{NP}$-hard.
[…]
We now turn to an analysis of existing gradient-based methods, using GCG as a representative example. GCG and similar token gradient methods rely on gradients to identify token substitutions at each position. For an input $x_{0}$, we compute the gradient of $f$ at $x_{0}$, denoted as $Df(x_{0})$. The gradient $Df(x_{0})$ has the same dimensionality as $x_{0}$. At position $j$, let $h=Df(x_{0})_{j}\in\mathbb{R}^{n}$ be the $j$-th component of $Df(x_{0})$. We can compute $k=\operatorname*{arg\,max}(h)$, which corresponds to the $k$-th token in the vocabulary $V$. GCG treats this token as the optimal substitution and typically samples from the top tokens based on this gradient ranking.
@@ 3.2 Hardness of Jailbreak Attack synthesis
The proof is presented in Appendix B. Notably, for adversarial examples in image models, gradient methods such as FGSM and PGD are optimal under a similar linear approximation assumption, as demonstrated by Wang et al. (2024). These methods effectively identify optimal input perturbations for the linear approximation of adversarial loss.
However, a key distinction lies in the nature of input perturbations. In image models, perturbations are restricted to small continuous $\ell_{p}$-balls, enabling accurate linear approximations. In contrast, token distances in language models can be substantial, reducing the accuracy of such approximations. As a result, applying token gradients to language models may be less effective.
@@ 3.3 Functional Homotopy method
The optimization problem $\min_{x\in\mathcal{X}}f(x)$ is a nonconvex and hard problem, whereas $\min_{x\in\mathcal{X}}g(x)$ is an easy optimization problem. As a result, $H(x,t)$ induces a series of easy-to-hard optimization problems.
One can then gradually solve this series of problems, by warm starting the optimization algorithm using the solution from the previous similar problem and eventually solve $\min_{x\in\mathcal{X}}f(x)$. Figure 2 illustrates an example of homotopy from $g(x)$ to $f(x)$. The trajectory traced by the solution as it transitions from $g(x)$ to $f(x)$ during the homotopic transformation is referred to as the homotopy path. Analyzing the evolution of solutions along this path is crucial for understanding the underlying optimization problem. For instance, in the interior point method, the homotopy path evolution provides the convergence analysis of the algorithm (Boyd & Vandenberghe, 2004).
@@
Constructing a homotopy offers various approaches. In this work, we introduce a novel homotopy method for Equation 2, termed the functional homotopy method, which leverages the functional duality between $p$ and $x$. Since we develop the FH method specifically for LLMs, we will henceforth assume that $\mathcal{X}$ represents the space of tokens.
@@
By allowing infinitesimal updates (learning rates), the gradient descent over the parameter space creates a homotopy between $F(p,x)$ and $F(p^{\prime},x)$, with $H(x,t=0)=F(p^{\prime},x)$ and $H(x,t=1)=F(p,x)$ for the homotopy method. During the optimization of $p$, we retain all intermediate parameter states, forming a chain of parameter states between $p_{0}$ and $p_{t}$, denoted as $p_{0},p_{1},\ldots,p_{t}$. Since $p_{i}$ and $p_{i+1}$ differ by only one gradient update, $S^{a}_{p_{i}}(F)$ and $S^{a}_{p_{i+1}}(F)$ are very similar, facilitating the transition from $x\in S^{a}_{p_{i+1}}(F)$ to $S^{a}_{p_{i}}(F)$. A formal description of the functional homotopy algorithm is provided in Algorithm 1. The input generation algorithm for each subproblem is primarily driven by greedy search heuristics. Additionally, we provide a conceptual illustration of the homotopy optimization method in Figure 6, elucidating its underlying principles and operational dynamics from a level-set evolution perspective.
@@
3.4 Application to Jailbreak Attack Synthesis
@@
For jailbreak attack generation, the objective is to persuade $M$ to provide an unaligned and potentially harmful response to a malicious query $x$ (e.g., “how to make a bomb?”), rather than refusing to answer. If $M$ is well-aligned, $T(M(p))$ should result in a refusal. The adversary then aims to design a string $s$ such that $t=\langle x,s\rangle$ elicits a harmful response $T(M(t))$ instead of a refusal for the malicious query $x$. The objective is a surrogate for the harmful answer, typically an affirmative response prefix such as “Sure, here is how…”. Zou et al. (2023); Liu et al. (2024a); Hu et al. (2024) have adopted similar formalizations for jailbreak generation.
@@ 4.1 Experimental Design
To quantify the similarity between these rankings, we employ the Rank Biased Overlap (RBO) metric (Webber et al., 2010). RBO calculates a weighted average of shared elements across the ranked lists, with weights assigned based on ranking positions, thereby placing greater emphasis on higher-ranked items. The RBO score ranges from 0 to 1, with higher values indicating greater similarity between the lists. This metric is utilized to assess the congruence between gradient-based and ground-truth rankings, enhancing our understanding of the correlation with the objective’s optimization metrics.
@@
We conduct a similar experiment to RQ2, but we record the number of search iterations used by each tool. Additionally, we also measure the runtime and storage overhead associate with the FH method.
@@ 4.2 Experimental Specifications
GR operates as a token-level search algorithm similar to GCG; however, it uses random selection for token substitutions rather than utilizing gradient information. This algorithm serves as an end-to-end implementation of 3 within Algorithm 1. Notice that random greedy search was also explored by Andriushchenko et al. (2024), as part of a bag of tricks applied in the work. Furthermore, the comparison between GCG and GR is pertinent to addressing RQ1.
@@
We use recent open source state-of-the-art models, in terms of performance and robustness. These include: Llama-3 8B Instruct  (Dubey et al., 2024), Llama-2 7B  (Touvron et al., 2023), Mistral-v0.3 7B Instruct  (Jiang et al., 2023) and Vicuna-v1.5 7B  (Chiang et al., 2023).
@@
For RQ1, we select 20 samples from the AdvBench dataset (Zou et al., 2023) and randomly choose four positions in the suffix for token substitution for each sample. For each query and position, we substitute all possible tokens (32 0003200032\,00032 000 for Llama-2, Mistral, and Vicuba; 128 256128256128\,256128 256 for Llama-3) and evaluate the jailbreak loss values using these inputs as ground truth, thereby establishing a ground truth ranking. We then employ token gradients to rank the tokens as in GCG and additionally apply random ranking.
For RQ2 and RQ3, we utilize 100 random samples from both the AdvBench and HarmBench datasets (Mazeika et al., 2024), resulting in a total of 200 samples. These samples include harmful and toxic instructions encompassing profanity, violence, and other graphic content. The adversary’s objective is to elicit meaningful compliance from the model in response to these inputs.
@@
We utilize the Llama-2 13B model, as provided by Mazeika et al. (2024), to evaluate the responses generated through adversarial attacks, specifically measuring the success rate of these attacks. In the context of jailbreak attack synthesis, the primary objective is to pass the evaluation by the judge, which effectively corresponds to the set $S_{p}^{a}(F)$ in Algorithm 1.
@@
The initial step of our FH method involves updating $p$, which effectively corresponds to model fine-tuning. To optimize memory and disk efficiency while preserving all intermediate parameter states, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021) for updating $p$. Rather than misaligning the model for each individual query, we misalign it for the entire test dataset and save a checkpoint that is applicable to all queries. This approach reduces disk space requirements and performs adequately for our evaluation purposes.
@@
We include other experimental specifications in Appendix C.